# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ 5.2. The Einstein-Weyl condition
Using the discussion in § 4.1, recall that a conformal connection $\nabla^{[g]}$ defines the projective connection (4.2). Using structure equations, the line field $\ell$ and the integrable rank 2 distribution $\mathscr{V}=\langle\widehat{\mu}_{*}\partial_{\theta^{1}},\widehat{\mu}_{*}\partial_{\theta^{2}}\rangle=\widehat{\mu}_{*}\mathrm{Ker}\{\omega^{0},\omega^{1},\omega^{2}\}\subset T\mathcal{T}^{[g]}$ satisfy the bracket generating relation $[\ell,\mathscr{V}]=T\mathcal{T}^{[g]},$ or, equivalently,
$\mathrm{d}\omega^{i}\equiv\varepsilon\theta^{i}{\,{\wedge}\;}\omega^{0}\mod\{\omega^{1},\omega^{2}\}.$
As a result, by the correspondence space theory of projective structures, e.g. see [ČS09, Section 4.4.3], $\mathcal{T}^{[g]}$ is endowed with a so-called path geometry where the paths are the integral curves of $\ell$ and project to the geodesics of $\nabla^{[g]}.$ If $[g]$ is Lorentzian then by (2.2) the integral curves of $\ell$ are the unique lift of time-like geodesics of $\nabla^{[g]}$ to $\mathcal{T}^{[g]}.$
@@
A conformal connection $\nabla^{[g]}$ is Einstein-Weyl if and only if, in terms of the coframing (4.7), the complex Pfaffian system $\{\zeta^{1},\zeta^{2}\}$ is integrable on $\mathcal{T}^{[g]}$. Subsequently, the CR complex structure on the contact distribution $\mathscr{C}\subset T\mathcal{T}^{[g]}$ descends to the 4-dimensional space of (time-like) geodesics of $\nabla^{[g]},$ also referred to as its minitwistor space.
@@
By Lemma 5.5, the space of (time-like) geodesics of $\nabla^{[g]},$ denoted by $\mathcal{S},$ coincides with the space of integral fields of $\widehat{\mu}_{*}\langle\partial_{\omega^{0}}\rangle\subset\mathcal{T}^{[g]}.$ Thus, the set of 1-forms $(s^{*}\omega^{1},s^{*}\omega^{2},s^{*}\theta^{1},s^{*}\theta^{2})$, for any section $s\colon\mathcal{S}\to\mathcal{C}_{0},$ give an adapted coframe on $\mathcal{S}.$ It is clear that if the complex Pfaffian system $\{\zeta^{1},\zeta^{2}\}$ is integrable then the induced complex structure on $\mathcal{S}$ with holomorphic distribution $\mathrm{Ker}\{\zeta^{1},\zeta^{2}\}\subset\mathbb{C}\otimes T\mathcal{S}$ is the descent of the CR holomorphic distribution given in the proof of Proposition 2.11 as $\mathrm{Ker}\{\omega^{0},\zeta^{1},\zeta^{2}\}.$
@@
for a conformal connection implies the Einstein-Weyl. It is clear that if $\nabla^{[g]}$ is Einstein-Weyl then (5.9) holds. Conversely, one needs to exploit the group action of $\mathrm{O}(p,q)$ on $\mathsf{P}_{ij},$ which at the infinitesimal level are encoded in the following Bianchi identities:
@@
mod $\{\omega^{0},\zeta^{1},\zeta^{2},\overline{\zeta}^{1},\theta^{3},\theta^{0}\}.$ Solving (5.9) and its differential consequences (5.10), one arrives at $\mathsf{P}_{(ij)}=\mathsf{P}_{00}\varepsilon_{ij}.$ Since the symmetric parts of $\mathsf{P}_{ij}$ and the Ricci tensor $R_{ij}$ are related as in Remark 4.2, it follows that $\nabla^{[g]}$ is Einstein-Weyl. ∎
@@
The existence of a complex structure on the minitwistor space Einstein-Weyl 3-manifolds was observed in [Hit82]. In relation to integrable $\mathrm{GL}(2,\mathbb{R})$-structures defined by scalar ODEs of odd order, a similar CR characterization via their twistor bundle is expected to hold. The twistor bundle of scalar ODEs of higher order has been studied in [Kry21] when the order is even.
@@ Acknowledgments
I would like to thank Daniel J. F. Fox and David Sykes for helpful conversations. I acknowledge partial support by the grant PID2020-116126GB-I00 provided via the Spanish Ministerio de Ciencia e Innovación MCIN/ AEI /10.13039/50110001103 as well as partial funding from the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and the UiT Aurora project MASCOT. Some parts of this article was done when I was at Masaryk University for which I would like to acknowledge the grant GA22-00091S. The EDS calculations are done using Jeanne Clelland’s Cartan package in Maple.

### v3

@@ 5.2. The Einstein-Weyl condition
Using the discussion in § 4.1, recall that a conformal connection $\nabla^{[g]}$ defines the projective connection (4.2). By structure equations, the line field $\ell$ and the integrable rank 2 distribution $\mathscr{V}=\langle\widehat{\mu}_{*}\partial_{\theta^{1}},\widehat{\mu}_{*}\partial_{\theta^{2}}\rangle=\widehat{\mu}_{*}\mathrm{Ker}\{\omega^{0},\omega^{1},\omega^{2}\}\subset T\mathcal{T}^{[g]}$ satisfy the bracket generating relation $\ell,\mathscr{V}]=T\mathcal{T}^{[g]},$ or, equivalently,
$\mathrm{d}\omega^{i}\equiv\varepsilon\theta^{i}{\,{\wedge}\;}\omega^{0}\mod\{\omega^{1},\omega^{2}\},$
which defines the multi-contact structure on $\mathbb{P}TM$ mentioned in our discussion in § 3.1 on the relation between geodesics of a projective structure and the integral curves of its geodesic spray, $\ell.$ In the correspondence space theory of projective structures, e.g. see [ČS09, Section 4.4.3], $\mathcal{T}^{[g]}$ is said to have a so-called path geometry where the paths are the integral curves of $\ell$ and project to the geodesics of $\nabla^{[g]}.$ If $[g]$ is Lorentzian then, by (2.2), the integral curves of $\ell$ are the unique lift of time-like geodesics of $\nabla^{[g]}$ to $\mathcal{T}^{[g]}.$
@@
A conformal connection, $\nabla^{[g]},$ is Einstein-Weyl if and only if, in terms of the coframing (4.7), the complex Pfaffian system $\{\zeta^{1},\zeta^{2}\}$ is integrable on $\mathcal{T}^{[g]}$. Subsequently, the CR complex structure on the contact distribution $\mathscr{C}\subset T\mathcal{T}^{[g]}$ descends to the 4-dimensional space of (time-like) geodesics of $\nabla^{[g]},$ also referred to as its minitwistor space.
@@
By Lemma 5.5, the space of (time-like) geodesics of $\nabla^{[g]},$ denoted by $\mathcal{S},$ coincides with the space of integral fields of $\widehat{\mu}_{*}\langle\partial_{\omega^{0}}\rangle\subset\mathcal{T}^{[g]}.$ Thus, the set of 1-forms $(s^{*}\omega^{1},s^{*}\omega^{2},s^{*}\theta^{1},s^{*}\theta^{2})$, for any section $s\colon\mathcal{S}\to\mathcal{C}_{0},$ gives an adapted coframe on $\mathcal{S}.$ It is clear that if the complex Pfaffian system $\{\zeta^{1},\zeta^{2}\}$ is integrable then the induced complex structure on $\mathcal{S}$ with holomorphic distribution $\mathrm{Ker}\{\zeta^{1},\zeta^{2}\}\subset\mathbb{C}\otimes T\mathcal{S}$ is the descent of the CR holomorphic distribution given in the proof of Proposition 2.11 as $\mathrm{Ker}\{\omega^{0},\zeta^{1},\zeta^{2}\}.$
@@
for a conformal connection implies the Einstein-Weyl condition. It is clear that if $\nabla^{[g]}$ is Einstein-Weyl then (5.9) holds. Conversely, one needs to exploit the group action of $\mathrm{O}(p,q)$ on $\mathsf{P}_{ij},$ which at the infinitesimal level are encoded in the following Bianchi identities:
@@
mod $\{\omega^{0},\zeta^{1},\zeta^{2},\overline{\zeta}^{1},\theta^{3},\theta^{0}\}.$ Solving (5.9) and its differential consequences (5.10), one arrives at $\mathsf{P}_{(ij)}=\mathsf{P}_{00}\varepsilon_{ij}.$ Since the symmetric parts of $\mathsf{P}_{ij}$ and the Ricci tensor $R_{ij}$ are related, as given in Remark 4.2, it follows that $\nabla^{[g]}$ is Einstein-Weyl. ∎
@@
The existence of a complex structure on the minitwistor space of Einstein-Weyl 3-manifolds was observed in [Hit82]. In relation to integrable $\mathrm{GL}(2,\mathbb{R})$-structures defined by scalar ODEs of odd order, a similar CR characterization via their twistor bundle is expected to hold. The twistor bundle of scalar ODEs of higher order has been studied in [Kry21] when the order is even.
@@ Acknowledgments
I would like to thank Daniel J. F. Fox and David Sykes for helpful conversations. I also thank the anonymous referee for many helpful and clarifying comments and corrections. I acknowledge partial support by the grant PID2020-116126GB-I00 provided via the Spanish Ministerio de Ciencia e Innovación MCIN/ AEI /10.13039/50110001103 as well as partial funding from the Tromsø Research Foundation (project “Pure Mathematics in Norway”) and the UiT Aurora project MASCOT. Some parts of this article was done when I was at Masaryk University for which I would like to acknowledge the grant GA22-00091S. The EDS calculations are done using Jeanne Clelland’s Cartan package in Maple.